OpenStudy (anonymous):

A triangle has area 49.5 cm squared. If the base of the triangle is 9 cm, what is the height of the triangle?

OpenStudy (anonymous):

49.5 = 9*.5*h 99= 9* 11=h

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem